40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 1 10:23:00 2009, steve, Update, PSL, MC2 damping restored again mc2again.jpg
    Reply  Thu Oct 1 11:39:47 2009, Jenne, Update, SUS, MC2 damping restored again 
       Reply  Thu Oct 1 14:22:28 2009, rob, Update, SUS, all suspensions undamped 
          Reply  Thu Oct 1 19:18:24 2009, Koji, Update, SUS, all suspensions undamped 
Message ID: 2034     Entry time: Thu Oct 1 11:39:47 2009     In reply to: 2033     Reply to this: 2036
Author: Jenne 
Type: Update 
Category: SUS 
Subject: MC2 damping restored again 

Quote:

 The EQ did not change the input beam pointing. All back to normal, except MC2 wachdogs tripped again.

 Round 3 for the day of MC2 watchdogs tripping.

ELOG V3.1.3-